Colony Inspection 26th May

Observations:

Weather had been good, temperatures above average at around 16C. Have been some heavy showers.
Chestnut still flowering and hawthorn now finishing.

Hive Record
Colony
Q
BIAS
Brood
Stores
CommentNext Inspection
AquaYY8 frames of brood, lot of play cupsabout 1 Kg in super.not much spacecheck for swarming and consider splitting
BlueY6 frames of brood, so growing quicklystarting to work superexpanding nicelycheck if still expanding
CyanNY4 frames BIAS with nice expanded patternstill consolidating into one brood boxbees a bit aggressivefind queen as she's not been seen for a while
EmeraldYY4 frames BIASOKno real change from last week.check if doing better, or unite?
FlaxNYjust 3 frames BIAS but also 3 charged QCs (took down)OKsurprised to see swarming behaviourcheck if still swaring and possibly unite to E